OpenStudy (anonymous):

All digits of a positive number M are odd and no two are the same. All digits of the positive integer N are even and no two are the same. If N is a multiple of M, determine the largest possible value of M.

OpenStudy (asnaseer):

largest M with those conditions must be 97531. largest N would be 86420 (as 0 is considered even). so largest NM = 86420 * 97531 = 8428629020 (I assume you wanted to know largest NM rather than largest M)

OpenStudy (anonymous):

I believe I want to know when M/N

OpenStudy (asnaseer):

in which case you want the largest M divided by the smallest N. smallest N would be 2 so largest M/N = 97531 / 2 = 48765.5

OpenStudy (anonymous):

I think it has to go along the line of something like odd integers is 2M+1 and even intergers are 2N

OpenStudy (asnaseer):

sorry - I didn't read the "If N is a multiple of M" - let me try again :-)

OpenStudy (anonymous):

ok no problem!

OpenStudy (asnaseer):

ok, since N is a multiple of M, then M must be less than or equal to the (largest N) / (smallest multiple) = 86420 / 2 = 43210 so M < 43210, and using the conditions given: largest M = 39751

OpenStudy (asnaseer):

correction, this doesn't satisfy the condition that N is a multiple of M with unique even digits. M = 7531 is the largest M that yields N = 8 * M = 60248